FIRST ANALYSIS OF THE ν_{5} BAND OF DNO_{3} (DEUTERATED NITRIC ACID) IN THE 11 μm REGION

J.KOUBEK, Institute of Chemical Technology, Department of Analytical Chemistry, Technická 5,166 28, Praha 6, Czech Republic, and Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France; \underline{A.PERRIN}, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du General de Gaulle, 94010 Créteil Cedex France; H. BECKERS, H. WILLNER, Anorg. Chemistry, University of Wuppertal, D-42119 Wuppertal, GERMANY.

The infrared spectrum of DNO_{3} (deuterated nitric acid) was recorded at high resolution (0.0018 cm ^{−1}) in the 800 -1400 cm ^{−1} region on the Bruker IFR 120 HR Fourier transform spectrometer of Wuppertal. The analysis of the ν_{5} A-type band of DNO_{3} centered at 887.656 cm ^{−1} was performed taking use of the ground state parameters achieved by Drouin et al. ^{a}. This band is perturbed by several resonances. ^{b}

aDrouin, Miller, Fry, Petkie, Helminger, Medvedev, J. Mol. Spectrosc. 236 (2006) 2934 bAcknowledgment: The work was supported through the Ministry of Education, Youth and Sports of the Czech Republic (research program LC06071) and scholarship of the French government (Bourse du Gouvernement Francais Bourse de Doctorat en co-tutelle).